OpenStudy (anonymous):

Jorge is selling raffle tickets for the carnival. He has sold 47 tickets and has 118 left to sell. How many tickets did he start with? A. t – 47 = 118 t = 165 165 tickets B. t + 47 = 118 t = 71 71 tickets C. t ÷ 2 = 47 t = 94 94 tickets D. t ÷ 2 = 118 t = 236 236 tickets

OpenStudy (anonymous):

@RoseDryer

OpenStudy (anonymous):

He started with 165 tickets

OpenStudy (anonymous):

Post the next ones. Post all the ones you've already sent too

OpenStudy (anonymous):

Which equation can be used to find the answer? A contractor needs 36 bags of cement to do a job. She already has 17 bags. How many more bags does she need to do the job? A. 2 · b = 36 b = 18 18 bags B. b + 17 = 36 b = 19 19 bags C. b ÷ 2 = 17 b = 34 34 bags D. b – 17 = 36 b = 53 53 bags

OpenStudy (anonymous):

Which equation can be used to find the answer? A state highway department plans to resurface 136 miles of highway during a six-month period. They have resurfaced 59 miles so far. How many more miles do they have left to resurface? A. m ÷ 2 = 59 m = 118 118 mi B. m + 59 = 136 m = 77 77 mi C. m – 59 = 136 m = 195 195 mi D. m ÷ 2 = 139 m = 278 278 mi

OpenStudy (anonymous):

Dania marks off 18.7 m of a playing field for a sports festival. She needs to mark off a total of 29 m to finish the field. How many more meters does Dania have left to mark off? A. f = 29 ÷ 2 f = 14.5 14.5 m B. f – 18.7 = 29 f = 47.7 47.7 m C. f = 18.7 · 2 f = 37.4 37.4 m D. f + 18.7 = 29 f = 10.3 10.3 m

OpenStudy (anonymous):

After two months, Rachel has saved enough money to buy an MP3 player. During the second month, she saved $87 and she pays $199 for the MP3 player. How much did Rachel save during the first month? A. m = 87 • 2 m = 174 $174 B. 199 + 87 = m m = 286 $286 C. m + 87 = 199 m = 112 $112 D. m – 87 = 199 m = 287 $287

OpenStudy (anonymous):

that is all @RoseDryer

OpenStudy (anonymous):

19 bags 77 mi 10.3 m $112

OpenStudy (anonymous):

thanks medal for you

OpenStudy (anonymous):

What value of n makes the equation true? n – 29.4 = 13.9 A. 15.5 B. 16.5 C. 42.3 D. 43.3 new ones six of them @RoseDryer

OpenStudy (anonymous):

he perimeter of this shape is 82 cm. Irregular quadrilateral What is the length of side d? cm

OpenStudy (anonymous):

43.3

OpenStudy (anonymous):

The length of a rectangle is three times the width. What is the perimeter of the rectangle if the width is 17 inches? A. 51 inches B. 68 inches C. 136 inches D. 204 inches

OpenStudy (anonymous):

Unit 2 | Lesson 6 Mid-Unit Checkpoint: Distance: Addition Equations Question Navigator Enter the answer. 4. How many feet are there in 456 in.? ft Question Resources

OpenStudy (anonymous):

In which measurement situations would the most appropriate unit of measurement be meters? Choose exactly two answers that are correct. A. width of a finger B. height of a tree C. depth of a well D. distance from the earth to the sun

OpenStudy (anonymous):

A rectangle has a length of 10 cm and a perimeter of 46 cm. Which equation could be used to find the width of the rectangle? A. 10w = 46 B. w + 10 = 46 C. 2w + 20 = 46 D. 2w + 10 = 46

OpenStudy (anonymous):

that is all

OpenStudy (anonymous):

 3w+3w+w+w 8w 8(17) 136 inches